OpenStudy (anonymous):

How would you convert 2i * sqrt20 (originally: i*sqrt80) into a+bi form?

jimthompson5910 (jim_thompson5910):

2i * sqrt20 = 2*2*sqrt(5)*i = 4*sqrt(5)*i So it's now in a+bi form where a=0 and b=4*sqrt(5)

OpenStudy (anonymous):

Cool, thank you for the reply.